The area of a trapezoid is 465ft what is its top base if the altitude is 28ft and the bottom base is 20ft. please solve it step by step.
Answer by nerdybill(7384)

If a and b are the lengths of the two parallel bases of a trapezoid, and h is its height, the area of the trapezoid is
area = 1/2 × h × (a + b)
.
plugging in what we know:
465 = 1/2 × 28 × (a + 20)
solving for 'a' -- the 'top of the trapezoid'
.
465 = 1/2 × 28 × (a + 20)
465 = 14 × (a + 20)
465/14 = (a + 20)
465/14 - 20 = a
33.214 - 20 = a
13.214 feet = a